OpenStudy (anonymous):

Determine whether each sequence could be geometric. what's the common ratio? 1.) 3, -6, 12, -24,.... 2.) -1, 5, -25, 125,....

hartnn (hartnn):

sequence will be geometric when the ratio of next term to previous term will be constant throughout ... so, for 1.) is -6/3 = 12/-6 = -24/12 =.....

hartnn (hartnn):

i mean are all those ratios equal ?

OpenStudy (dan815):

that was alll wrong

OpenStudy (dan815):

again!

OpenStudy (anonymous):

lol

OpenStudy (dan815):

lmao!

OpenStudy (dan815):

ratios are basically 2 and 5 but (-1)^n in front of it

OpenStudy (anonymous):

for 1 or 2?

OpenStudy (dan815):

1)= 2 and 2) = 5

OpenStudy (dan815):

but where does that -1 go

OpenStudy (dan815):

lemme try to write this formula again >_>

OpenStudy (anonymous):

lol okay!(:

OpenStudy (dan815):